Jessica has a rectangular flag that is 18 inches long and 11 inches wide. What is the perimeter of the flag? Answer in units of inches.
Answer by fractalier(6550)

Perimeter P = 2L + 2W
Here that would be
P = 2(18) + 2(11) = 36 + 22 = 58 inches.